# Automatic Fused Multimodal Deep Learning for Plant Identification

### v1

@@ 2.4.1 Unimodal Model Training
As mentioned in Section 2.2.1, we train only a classification layer of each unimodal model. This results in 551,612 trainable parameters per model. Each model is trained with identical parameters. We apply an Adam optimizer, initial learning rate of 0.001, utilizing exponential decay of 0.95 every 300 steps during the training of unimodal models. Moreover, models undergo training for 1000 epochs with a batch size of 128 and shuffling. However, we incorporate early stopping with the patience of 10 and the best weight restoration.
To minimize the impact of class imbalance, we employ a weighted cross-entropy loss, where a weight for each class is computed according to the formula:
$w_{c}=1-\frac{N_{c}}{N}$ (2)
where $N_{c}$ is the number of instances of a class $c$ within the training data, and $N$ is a total number of instances. Additionally, the L1 and L2 regularizations of $10^{-5}$ are applied. Moreover, label smoothing of 0.1 is applied to the loss.
@@ 2.4.2 Multimodal Fusion Architecture Search Execution
During the MFAS algorithm procedure, each multimodal architecture is trained with Adam optimizer, with a learning rate of 0.001 and a weighted cross-entropy loss, with class weights determined according to Equation 2. We fix the number of neurons per fusion layer to 64. Each architecture undergoes training for 2 epochs with a batch size of 128 and shuffling. Given the imbalanced nature of the dataset, we monitor the training progress by checking the F1 macro metric on the validation set. Similar to Perez-Rua et al. (2019), we initialize our temperature scheduler with $t_{max}=10$, $t_{min}=0.2$, and temperature decay $d=4$.
To make models robust to missing modalities (as discussed in Section 2.3), we adopt the multimodal dropout technique proposed by Cheerla and Gevaert (2019). Inspired by a regular dropout, multimodal dropout involves dropping entire modalities during training. This encourages the model to build representations robust to missing modalities. Each modality of a sample is dropped with a similar probability. We utilize a low multimodal dropout rate of 0.125 because many classes are completely devoid of certain modalities.
In this study, we execute the MFAS algorithm for 4 iterations, with 3 progression levels each and set the number of sampled architectures to 50. It is crucial to note that exploring all 1250 initial architectures at the start of the algorithm is time-consuming, so they have been explored in 20 parallel batches. Subsequently, based on the scores obtained, the algorithm has been restarted with the 50 best architectures used as initial ones.
@@ 2.4.3 Final Model Training
After the execution of the MFAS algorithm, we select the 10 most performant architectures and train for 100 epochs. Early stopping is employed with a patience of 10, and exponential learning rate decay of 0.95 is applied each epoch. The other hyperparameters remain the same as during the search to determine the best architecture.
Subsequently, we tune the hyperparameters of the best architecture carefully. Initially, the optimal number of neurons per layer is investigated. As suggested by Perez-Rua et al. (2019), small weight matrices are utilized during the algorithm to enhance its speed and reduce memory consumption. However, when focusing on a single architecture, this limitation is no longer applicable. Next, dropouts, various regularization techniques, and other hyperparameters are tested to find an optimal configuration. Finally, the resulting model is trained to completion for 100 epochs with early stopping with a patience of 10. At this stage, the unimodal backbones within it are unfrozen and fine-tuned with a lower learning rate to potentially further enhance performance.
@@ 2.5.1 Establishing Baseline
We implement the late fusion using the averaging strategy, similar to Baltrusaitis et al. (2019). In this approach, the final prediction for an instance $x$ is calculated according to the formula:
@@ 2.5.2 Comparison with the Baseline
For each class, we compute both micro and macro averages of each metric, except for accuracy. In micro averaging, all $TP$, $TN$, $FP$, and $FN$ values are accumulated before calculating the metric, whereas in macro averaging, a metric is first calculated for each class and then averaged across all classes. The discrepancy between these values provides insight into the performance variation across classes. In conjunction with these metrics, top-5 and top-10 accuracies are also calculated.
@@ 2.5.3 Robustness to Missing Modalities
To enable a comparison between unimodal models and multimodal models, we also collect the same metrics discussed in Section 2.5.2 for individual unimodal models. Moreover, to facilitate an understanding of the robustness capabilities of the proposed model in the absence of certain modalities, we collect the mentioned metrics for the proposed model and the baseline on different subsets of modalities.
@@ 3.1 Performance of Unimodal Models
Accuracy 0.6411 0.5147 0.6759 0.4784
Top-5 accuracy 0.8281 0.7178 0.8416 0.7013
Top-10 accuracy 0.8844 0.7863 0.8895 0.7749
Precisionmicro 0.6411 0.5147 0.6759 0.4784
Precisionmacro 0.6244 0.4532 0.6883 0.4184
Recallmicro 0.6411 0.5147 0.6759 0.4784
Recallmacro 0.5574 0.3815 0.6248 0.3888
F1micro 0.6411 0.5147 0.6759 0.4784
F1macro 0.5671 0.3916 0.6276 0.3814
Table 1: Performance Metrics of Unimodal Models
According to Tabe 1: (i) It is evident that the fruit and flower modalities exhibit higher scores across all metrics compared to leaves and especially stems. This is in line with expectations, considering that stems are the least discriminative organs (Nhan et al., 2020). (ii) Micro metrics demonstrate higher values compared to macro ones. This discrepancy should be attributed to differences in performance among classes. As mentioned in Section 2.3, the dataset is imbalanced, which has affected the overall performance. (iii) Higher macro precision values compared to macro recall suggest that, on average, the models tend to miss correct instances rather than misclassify them. (iv) Relatively high top-$N$ accuracies indicate that the models are typically close to identifying the correct class.

### v2

@@ 2.4.1 Unimodal Model Training
Each base MobileNetV3Small model is appended with an intermediate dense layer (with batch normalization) and a classifier layer. Initially, we train only these appended layers using the default Adam optimizer with an exponentially decaying LR and early stopping (ES), after which the best weights are restored based on validation loss. Subsequently, we fine-tune from the top layer down to and including the 4th inverted residual block (IRB) for the leaf and fruit models, and all layers for the flower model. Fine-tuning did not yield performance improvements for the stem model, so it is omitted there. During fine-tuning, ES is delayed until the 40th epoch. Table 2 provides the complete list of unimodal model hyperparameters.
Models: Flower Leaf Fruit Stem
Optimizer Adam Adam Adam Adam
Initial LR $10^{-3}$ $5^{-4}$ $5^{-4}$ $10^{-3}$
LR decay rate $0.95$ $0.95$ $0.95$ $0.95$
LR decay steps 200 200 200 200
Batch size 256 256 256 256
Epochs 1000 1000 1000 1000
ES patience 10 10 10 10
Classifier L1 $10^{-5}$ 00 00 00
Classifier L2 $10^{-5}$ $10^{-3}$ $10^{-3}$ $10^{-2}$
Classifier dropout 0.4 0.1 0.2 0.0
Intermediate neurons 1024 1024 512 1024
Intermediate layer L2 0 $10^{-5}$ 0 0
Fine-tuning:
Fine-tuning True True True False
Initial LR $10^{-4}$ $10^{-4}$ $8^{-5}$
ES from epoch 40 40 40
Fine-tuned layers All Up to 4th IRB Up to 4th IRB
Retraining on combined training and validation sets:
ES False False False False
Epochs 160 140 120 110
Fine-tuning epochs 160 120 180
Hyperparameters remain unchanged during fine-tuning or retraining on combined
datasets unless specified otherwise.
Table 2: Hyperparameters of Unimodal Models
To minimize the impact of class imbalance, we employ a weighted cross-entropy loss, where a weight for each class $c\in C$ is computed according to the formula:
$w_{c}=\frac{N}{|C|}\times N_{c}$ (3)
where $N_{c}$ is the number of instances of a class $c$ within the training data, and $N$ is a total number of instances.
Additionally, we shuffle the data at each epoch and randomly apply image augmentations that have shown to enhance performance, as outlined in Table 3.
Models: Flower Leaf Fruit Stem
Contrast True False True False
Horizontal flip True True True True
Vertical flip False True False False
Contrast adjustment computes a mean $\mu$ for each image channel $c$ and
sets it for each pixel as $(c-\mu)\times k+\mu$, where $k\sim U(0.75,1.25)$.
Table 3: Image Augmentations for Unimodal Models
To fully utilize the available data, we train two versions of each model: one on the training set only, and another on the combined training and validation sets. The first model version is used during MFAS execution (Section 2.4.2) and while tuning the final model (Section 2.4.3). Once hyperparameter tuning is complete, the validation set is no longer required, so we merge it with the training set and retrain the models on this combined data for evaluation purposes. Without a validation set, we cannot apply ES based on validation performance; therefore, we determine an appropriate number of training epochs based on the first model version, increasing it slightly to account for the additional data. The number of epochs used for retraining on the combined sets is provided in Table 2.
@@ 2.4.2 Multimodal Fusion Architecture Search Execution
In the MFAS algorithm procedure, each multimodal architecture is trained using the Adam optimizer with a learning rate of 0.001 and a weighted cross-entropy loss, where class weights are calculated according to Equation 3. The number of neurons per fusion layer is set to 64. Each architecture is trained for 2 epochs with a batch size of 256. As batching significantly impacts training speed, we cache batches and shuffle them each epoch in buffers of 12, rather than creating new random batches each time. Given the limited number of epochs per architecture during the algorithm’s execution, this approach offers a favorable trade-off between maintaining batch randomness and optimizing training speed. Given the imbalanced nature of the dataset, we score the architectures using the F1macro metric on the validation set. Similar to Perez-Rua et al. (2019), we initialize our temperature scheduler with $t_{max}=10$, $t_{min}=0.2$, and temperature decay $d=4$.
In this study, we execute the MFAS algorithm for 5 iterations, with 4 progression levels each and set the number of sampled architectures to 50. It is important to note that exploring all 2592 initial architectures at the start of the algorithm is time-consuming; therefore, they were evaluated in 20 parallel batches. Subsequently, the algorithm was restarted with the obtained results and a pre-trained surrogate, continuing from the second progression level with the 50 best architectures sampled from the first level.
@@ 2.4.3 Final Model Training
To identify the final architecture after executing the MFAS algorithm, we select the 10 most performant configurations and train each for 100 epochs. We employ ES with a patience of 10 and an initial LR of 0.001, which decays exponentially, reaching 95% of its value every 200 steps. We also apply batch normalization after each fusion layer, while retaining the other hyperparameters used during the architecture search.
Subsequently, we tune the hyperparameters of the best architecture carefully. Initially, the optimal number of neurons per layer is investigated. As suggested by Perez-Rua et al. (2019), small weight matrices are utilized during the algorithm to enhance its speed and reduce memory consumption. However, when focusing on a single architecture, this limitation is no longer applicable. Next, we test various configurations of LR, its decay, dropout rates, and other regularization techniques to identify the optimal setup. Additionally, we explore fine-tuning this model to potentially enhance its performance further.
After hyperparameter tuning, we combine the training and validation sets, as detailed in Section 2.4.1, and retrain the model. We also use the corresponding versions of the unimodal models trained on the merged sets. To enhance the model’s robustness to missing modalities (as discussed in Section 2.3), we implement the multimodal dropout (MD) technique proposed by Cheerla and Gevaert (2019). Inspired by a regular dropout, MD involves dropping entire modalities during training. In our case, this means replacing pixel values of an organ image with zeros. This encourages the model to build representations robust to missing modalities. Each modality of a sample is dropped with a similar probability. We utilize a low dropout rate of 0.125 because many classes are completely devoid of certain modalities. Given the challenge of validating this technique during training, we train two versions of the final model: one with MD and one without, and subsequently evaluate both.
@@ 2.5.1 Establishing Baseline
We implement the late fusion using the averaging strategy (Baltrusaitis et al., 2019). In this approach, the final prediction for an instance $x$ is calculated according to the formula:
@@ 2.5.2 Comparison with the Baseline
For each class, we compute macro averages of each metric, excluding accuracy. We omit micro averaging since it accumulates all $TP$, $TN$, $FP$, and $FN$ values before calculating the metric, resulting in a value equivalent to accuracy. In contrast, macro averaging calculates each metric for individual classes and then averages these values across all classes. The difference between accuracy and macro-averaged metrics provides insight into the performance variation across classes. In conjunction with these metrics, top-5 and top-10 accuracies are also calculated.
@@ 2.5.3 Robustness to Missing Modalities
To enable a comparison between unimodal models and multimodal models, we also collect the same metrics discussed in Section 2.5.2 for individual unimodal models. Moreover, to facilitate an understanding of the robustness capabilities of the proposed model in the absence of certain modalities, we collect the mentioned metrics for the proposed model and the baseline on different subsets of modalities. To evaluate the statistical significance of the differences between these metrics, we apply McNemar’s test.
@@ 3.1 Performance of Unimodal Models
Accuracy 0.6796 0.4294 0.4936 0.2698
Top-5 accuracy 0.8597 0.6508 0.7129 0.4622
Top-10 accuracy 0.9082 0.7293 0.7773 0.5652
Precisionmacro 0.5998 0.2304 0.3215 0.1298
Recallmacro 0.5794 0.2334 0.3444 0.1449
F1macro 0.5702 0.2181 0.3143 0.1247
Table 4: Performance Metrics for Unimodal Models
According to Table 4: (i) It is evident that the fruit and flower modalities exhibit higher scores across all metrics compared to leaves and especially stems. This is in line with expectations, considering that stems are the least discriminative organs (Nhan et al., 2020). (ii) Accuracy demonstrates higher values compared to macro metrics. This discrepancy should be attributed to differences in performance among classes. As mentioned in Section 2.3, the dataset is imbalanced, which has affected the overall performance. (iii) Despite that, similar yet lower precision and recall values suggest that models’ performance is consistent on more common classes but poorer on less common. (iv) Relatively high top-$N$ accuracies indicate that the models are typically close to identifying the correct class.